Hydrated copper sulfate can be represented as CuSO4• H2O. When 10.00g of the hydrated salt is heated in an oven to drive off the water of crystallisation, the mass of anhydrous (without water) salt remaining is 6.39g. What is the formula of the hydrated salt?
1
Expert's answer
2020-05-07T14:08:26-0400

"CuSO_4*NH_2O"

mass of salt =10,00g

mass of dry salt = 6,39g

mass of water = 3,61g

"n=m\/M" , n - chemical amount of substance; m - mass of substance; M - molar mass of substance. 

n(H2O)=3,61g/18g/mol=0,2mol

n(CuSO4)=6,39g/160g/mol=0,04mol

N=n(H2O)/n(CuSO4)=0,2/0,04=5

Answer:CuSO4*5H2O